Question 1.7 [2, 2, 2, 2]


If A and B are independent events with P A( ) 0.52 = and P B( ) 0.21 = , find the following:


a) PA B ( ) ∪


b) PA B ( ) ∩


c) PA B ( ) ∪


d) Are A and B disjoint events? Motivate your answer!


Question 1.8 [3]


A certain washing machine factory has found that 15% of its washing machines manufactured in the


factory break down and are returned in the first year of operation. Suppose that 32 machines are


purchased by a laundromat from this washing machine factory, find the probability that at least one


washing machine breaks down in the first year of operation?

1
Expert's answer
2022-06-06T14:45:22-0400

1.7.

"P(A\cap B)=P(A)P(B)=0.52(0.21)=0.1092"

a)


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=0.52+0.21-0.52(0.21)=0.6208"

b)


"P(A^C\cap B^C)=1-P(A\cup B)"

"=1-0.6208=0.3792"

c)


"P(A^C\cup B^C)=P(A^C)+P(B^C)-P(A^C\cap B^C)"

"=(1-0.52)+(1-0.21)-0.3792=0.8908"




"P(A^C\cup B^C)=1-P(A\cap B)"

"=1-0.1092=0.8908"

d)

If two events are disjoint, then the probability of them both occurring at the same time is 0.

Since "P(A\cap B)=0.1092\not=0," then "A" and "B" are not disjoint events.


1.8.

Let "X=" the number of washing machine which breaks down in the first year of operation: "X\sim Bin (n, p)."

Given "n=32,p=0.15, q=1-p=0.85."


"P(X\ge 1)=1-P(X=0)"

"=1-\dbinom{32}{0}(0.15)^{0}(0.85)^{32-0}"

"=0.9944867762"